suppose A is 5x5 matrix with eigenvector

5 0 5 -15

and an eigenvalue lambda= -1.

The third row of A is

[1 0 -1 k 20]what is k?
Answer #1

Solution:

Ax=\lambda x

where A is a matrix, x is eigenvector and lambda is an eigenvalue

\\\begin{bmatrix} 1&0&-1&k&20 \end{bmatrix}\begin{bmatrix} 5\\ 0\\ -5\\ -15\\ 1 \end{bmatrix}=(-1)\begin{bmatrix} 5\\ 0\\ -5\\ -15\\ 1 \end{bmatrix}=\begin{bmatrix} -5\\ 0\\ 5\\ 15\\ -1 \end{bmatrix}

The third row of A is given

[5+5-15k+20]=5

\\30-15k=5\\15k=25\\3k=5\\\\k=\frac{5}{3}